# envname-P envname#1

### v2

@@
Abstract.
In this paper, we introduce the novel task of Open-domain Urban Itinerary Planning (OUIP), a paradigm designed to generate personalized urban itineraries from user requests articulated in natural language. This approach is different from traditional itinerary planning, which often restricts the granularity of user inputs, thus hindering genuine personalization. To this end, we present ItiNera, an OUIP system that synergizes spatial optimization with large language models (LLMs) to provide services that customize urban itineraries based on users’ needs. Upon receiving the user’s itinerary request, the LLM first decomposes it into detailed components, identifying key requirements, including preferences and dislikes. Then, we use these specifics to select candidate POIs from a large-scale collection using embedding-based Preference-aware POI Retrieval. Finally, a preference score-based Cluster-aware Spatial Optimization module clusters, filters, and orders these POIs, followed by the LLM for detailed POI selection and organization to craft a personalized, spatially coherent itinerary. Moreover, we created an LLM-based pipeline to update and personalize a user-owned POI database. This ensures up-to-date POI information, supports itinerary planning, pre-trip research, POI collection, recommendations, and more. To the best of our knowledge, this study marks the first integration of LLMs to innovate itinerary planning, with potential extensions for various urban travel and exploration activities. Offline and online evaluations demonstrate the capacity of our system to deliver more responsive, personalized, and spatially coherent itineraries than current solutions. Our system, deployed on an online platform, has attracted thousands of users for their urban travel planning.
Itinerary Planning; Intelligent Transportation; Location-based Services; Spatial Optimizations; Large Language Models
1. Introduction
As a novel form of urban travel and exploration, city tours invite travelers to wander through city streets, delve into historical sites, and immerse themselves in local culture. Hence it offers a more efficient, dynamic, and cost-effective travel experience compared to traditional tourism. Importantly, COVID-19 shifted travel trends towards crowd-free, health-conscious activities deeply connected with local contexts, propelling city tours to global popularity (Germano, 2023). However, planning a city tour is essentially an urban itinerary planning problem (Halder et al., 2024), a complex procedure that involves gathering travel-related information, selecting POIs, mapping out routes, and customizing to accommodate diverse user needs. An enjoyable city tour typically requires a significant time commitment in preparation.
Existing itinerary planning studies focus on traditional tourism. They typically consider coarse-grained user requirements, such as must-see POIs (Taylor et al., 2018), categories(Bolzoni et al., 2014), geographical considerations (Rani et al., 2018, 2019), and time budget (Zhang and Tang, 2018; Hsueh and Huang, 2019), or a combination of them (Yochum et al., 2020) as constraints to improve the logical arrangement of sites and optimize the quality of an itinerary (Chen et al., 2013; Sylejmani et al., 2017). We refer to this approach as spatial optimization, and it can maintain the quality of POIs and spatial coherency of the planned itinerary. However, the spatial optimization approach exhibits a lack of personalization and fails to ensure overall quality, thereby failing to satisfy the demands of users. A high-quality itinerary for a city tour should transcend a mere list of POIs, and importantly, it calls for alignment with users’ specific requirements and lifestyles.
Figure 1. The OUIP problem and the deployed OUIP service.
To elaborate on this, we summarize the highlighting features of city tours that differentiate itself from traditional tourism:
  • Dynamic Information: Unlike conventional tourism, where POIs remain relatively static, city tours involve rapidly changing POIs. This requires up-to-date information and allows for the exploration of temporary events, pop-up venues, and other fleeting experiences that traditional itineraries might overlook.
  • Individually Tailored: The planning of a city tour places significant emphasis on personal tastes and interests. Unlike traditional tourism, which often focuses on universally recognized POIs, city tours prioritize the unique preferences of the individual, making each itinerary deeply personalized.
  • Fine-grained Exploration: City tours encourage a detailed exploration of urban environments and a deeper understanding and appreciation of the city’s culture, architecture, and history, which goes beyond the surface-level glimpse offered by traditional tourism.
  • Diverse Constraints: City tour planning takes into account more complex constraints, necessitating more intricate planning. This includes implicit constraints that align with a person’s lifestyle and explicit constraints such as specific traveler interests or accessibility requirements.
For example, an ideal itinerary might start at a coffee shop in the morning, traverse through selected preferred POIs without detours, and conclude at an affordable yet atmospheric restaurant. An itinerary of this granularity not only fulfills travelers’ distinct preferences but significantly enriches the travel experience.
One can see that it is challenging to directly apply conventional itinerary planning approaches due to the above-discussed features of city tours. Ideally, users would articulate their preferences in a single sentence and, in return, receive a personalized, spatially coherent, well-designed itinerary that satisfies their requirements. Based on these observations, there is a need for a new itinerary planning paradigm that can integrate detailed user preferences to craft personalized and fine-grained travel experiences. In view of these, we propose the new task of open-domain urban itinerary planning (OUIP) in Definition 1.
Definition 1 (OUIP-layman’s version).
Open-domain Urban Itinerary Planning (OUIP) aims to generate personalized urban travel itineraries based on user preferences in natural language.
The term “open-domain” involves generating responses to queries across a wide range of topics without domain-specific constraints. Similarly, our proposed OUIP problem aims to generate personalized itineraries by interpreting users’ preferences expressed in natural language, as depicted in Fig. 1.
Recently, LLMs have demonstrated impressive abilities in language comprehension and instruction-following, enabling them to understand user needs, handle simple planning tasks, and generate responses based on their knowledge (OpenAI, [n. d.]). Recent trends propose employing LLMs for itinerary planning problems (Xie et al., 2024). However, the limitations of LLMs become apparent in itinerary planning due to reliance on non-real-time information, incomplete knowledge, and a lack of spatial awareness. Consequently, they are unable to independently deliver a satisfactory user experience in itinerary planning. The itineraries generated directly by LLMs could be circuitous, lack of detail, and include hallucinated, non-existent POIs, rendering them less practical for actual use.
Beyond the methodological gaps, evaluating various approaches to OUIP presents its own set of difficulties. First of all, there is a general lack of datasets for evaluating the OUIP approaches, and the true city tour plans are usually presented in the format of unstructured texts and images. The baseline methods for OUIP are also limited. More critically, considering the open-domain nature of the OUIP task, it is insufficient to rely solely on traditional rule-based metrics for evaluation.
Given above challenges, to address the task of OUIP, this paper proposes ItiNera, a holistic OUIP system that synergizes spatial optimization with LLMs. Specifically, ItiNera features a User-owned POI Collection module capable of automatically scraping and extracting POI features from web content, dynamically updating a user-owned, personalized POI database to ensure both the customization and timeliness of POI information, laying the groundwork for subsequent OUIP tasks. ItiNera consists of five modules: User-owned POI Collection, Request Decomposition, Preference-aware POI Retrieval, Cluster-aware Spatial Optimization, and Itinerary Generation. All five modules are LLM-assisted, and details will be presented later. Importantly, as the OUIP is a novel task, we collect and build the validation dataset and develop a series of metrics for a comprehensive evaluation framework. Lastly, ItiNera has been deployed as a real-world service that attracts thousands of users within months.
Overall, the main contributions of our work are as follows:
  • We introduce the Open-domain Urban Itinerary Planning (OUIP) problem, designed to provide personalized urban travel itineraries based on users’ natural language inputs.
  • We develop an LLM-based OUIP system, ItiNera. To the best of our knowledge, this is the first study that combines spatial optimization with LLMs to create spatially coherent, fine-grained urban itineraries tailored to users’ requests.
  • We designed several metrics employing both rule-based assessments and GPT-based metrics to measure the quality and personalization of generated itineraries, which could enlighten future OUIP research by providing a comprehensive metric framework.
  • Extensive experiments on real-world dataset demonstrate that ItiNera delivers the best performance in creating personalized, spatially coherent, high-quality urban itineraries that meet user requirements; ItiNera is also deployed as the core system of TuTu to provide online OUIP service. User feedback validates the effectiveness of our system in real-world scenarios.
Figure 2. An overview of the ItiNera system.
2. Related Works
2.1. LLMs in Urban Application
Since ChatGPT (OpenAI, [n. d.]), Large Language Models (LLMs) have received widespread attention from both the academic and industrial communities. They have demonstrated strong capabilities in tasks such as language processing, instruction-following, planning, and learning. Existing applications of LLMs are concentrated in agent (Wang et al., 2023c; Xi et al., 2023) and prompt (Yao et al., 2023; Wei et al., 2022) design. However, the urban applications using LLMs remain in the early stages. Recent studies highlight the potential of LLMs in urban data processing (Yan et al., 2023), showing their impact on urban planning (Zhou et al., 2024). These works reveal LLMs’ capabilities in predicting human mobility patterns (Mo et al., 2023; Xue et al., 2022), including during public events (Liang et al., 2023), and emphasize their predictive strength (Xue and Salim, 2023; Wang et al., 2023a). In transportation, LLMs contribute to traffic safety analysis (Zheng et al., 2023a), enhance traffic forecasting (de Zarzà et al., 2023), and automate accident report generation (Zheng et al., 2023c), demonstrating their wide applicability in urban transportation. (Zheng et al., 2023b)
Leveraging LLMs for automated travel planning has recently gained public interest. Concurrently, TravelPlanner (Xie et al., 2024) proposes a sandbox environment with various tools for benchmarking LLMs on long-term (minimum 3 days) travel planning, revealing LLMs’ current limitations in handling complex planning tasks. Given this, we present ItiNera, a holistic system designed for OUIP. Unlike TravelPlanner, our system focuses on fine-grained OUIP, addressing urban itinerary planning within a single day, but can be seamlessly extended to multi-day travel planning.
2.2. Itinerary Planning
Current research on IP focuses on creating itineraries based on a set of POIs. Some methods directly optimize the spatial utilities of the itinerary to obtain spatially coherent itineraries, while others define IP as an Orienteering Problem (OP) and consider constraints that include time (Zhang and Tang, 2018; Hsueh and Huang, 2019), space (Rani et al., 2018, 2019), must-see POIs (Taylor et al., 2018), categories (Bolzoni et al., 2014), and their combinations (Gionis et al., 2014; Yochum et al., 2020), optimizing POI utilities such as location popularity (Varakantham and Kumar, 2013; Gunawan et al., 2014) to indirectly ensure the spatial coherence and quality of the itinerary. Although these methods can ensure the quality of POIs in the itinerary, their ability to personalize is limited. Some recommendation-based methods (Ho and Lim, 2022; Tang et al., 2022) could be applied to the IP task. However, these methods often depend on historical user behavior data. Overall, existing IP methods struggle with open-domain, user natural-language inputs, failing to generate personalized itineraries, making them unsuitable for OUIP.
3. Preliminaries
Definition 2 (User Request).
A user request $\small r$ refers to a text string that includes the user’s natural language description of needs for an itinerary. The user request potentially consists of both hard requirements (i.e., constraints) and general preferences for the itinerary.
Definition 3 (Point of Interest).
A Point of Interest (POI) $\small p$ represents a specific location, like landmarks, restaurants, or cultural sites, characterized by its geographical location, category, descriptions, and other relevant context features.
Definition 4 (OUIP-full version).
Given a user request $\small r$ and the user-owned POI set $\small\mathcal{P}=\{p_{j}\}_{j=1}^{N}$, the OUIP problem aims to find an itinerary generator $\small\mathcal{G}_{\Theta}$ to select and order a subset of POIs from $\small\mathcal{P}$ to create a coherent travel itinerary that optimally aligns with the user’s requests $\small r$ while adhering to spatio-temporal considerations:
(1) $\small I\sim\mathcal{G}_{\Theta}\left(\mathcal{P}|r\right),$
where $\small\Theta$ is the parameters of the itinerary generator $\small\mathcal{G}$, and the generated itinerary $\small I$ is an ordered list of POIs.
4. Methodologies
4.1. Overview
In this section, we provide details of ItiNera. As depicted in Fig. 2, we design five modules to enable an OUIP service. We first incorporate a User-owned POI Collection (UPC) module to facilitate travel information gathering and user-owned POI database construction. To plan an itinerary that meets a user’s request, we design a Request Decomposition (RD) module to accurately interpret and extract independent user preferences. Subsequently, we develop a embedding-based Preference-aware POI Retrieval (PPR) Module that fetches the top-K most relevant POIs to form the retrieved POI set for the itinerary. To ensure a spatially coherent itinerary, we employed a Cluster-aware Spatial Optimization (CSO) module to spatially filter and rank retrieved POIs by solving the hierarchical traveling salesman problem, thereby obtaining the candidate POIs. Lastly, the Itinerary Generation (IG) module integrated the ordered candidate POI set with multiple constraints to construct the prompt for LLM’s completion. This approach enables LLM to generate travel routes that are spatio-temporally rational and align with user requests. We provide details for each module in the following subsections.
4.2. User-owned POI Collection
In OUIP, the way of travel information acquisition has undergone a significant transformation from conventional travel planning due to technological advancements. Base on our user and market research, travelers increasingly rely on social media for pre-trip information gathering. Notably, user-generated content is more authentic and current in terms of customer experience than official sources of information. There are existing methods to gather pre-trip information, such as using LLMs to call tools for searching relevant POIs like autoGPT (Gravitas, 2023). However, these LLM-agent-based methods typically consume time and are not suitable for instant use. Moreover, they require users to have relatively specific targets or requirements, which hinder users to discover new POIs.
@@ 4.2. User-owned POI Collection
Details of this pipeline and the constructed database format are further explained in Appendix E and B, respectively.
4.3. Request Decomposition
Upon receiving user requests, we leverage the advanced natural language processing capabilities of LLMs to extract structured user requests in this module.
Based on our received requests, we’ve observed that a single user request $\small r$ can be decomposed into multiple independent subrequests, each could be generally classified according to their granularity, specificity, and attitude. For granularity, we have (1) POI-level subrequests and (2) itinerary-level subrequests. In terms of specificity, there are (1) specific subrequests and (2) vague subrequests. Regarding attitude, we identify (1) positive subrequests and (2) negative subrequests. Thus, we prompt LLM to decompose the user request $\small r$ base on the above identified categories. We formally use $\small\operatorname{LLM}_{\theta}$ to denote a pre-trained LLM with parameters $\small\theta$:
(2) $\small\mathcal{R}\sim\operatorname{LLM}_{\theta}\left(\,\mathbb{P}_{RD}\left(r\right)\right),$
where $\small\mathbb{P}_{RD}$ wraps the request $\small r$ with instructions and few-shot input-output examples, the result is a set of decomposed independent structured subrequests $\small\mathcal{R}=\{\mathbf{r}_{i}\}_{i=1}^{|\mathcal{R}|}$. Each subrequest $\small\mathbf{r}_{i}$ comprises four keys: ‘pos’, ‘neg’, ‘mustsee’, ‘type’, with their corresponding values $\small\mathbf{r}_{i}^{\operatorname{pos}}$, $\small\mathbf{r}_{i}^{\operatorname{neg}}$, $\small\mathbf{r}_{i}^{\operatorname{must}}$, $\small\mathbf{r}_{i}^{\operatorname{type}}$. Here, ‘pos’ and ‘neg’ are strings indicating the user’s likes and dislikes contained in the subrequest, ‘mustsee’ is a boolean (true or false), and ‘type’ can be one of [”start”, ”end”, ”POI”, ”itinerary”], indicating whether the subrequest targets the starting POI, ending POI, a specific POI, or the entire itinerary, respectively. To explain each field in a subrequest $\small\mathbf{r}_{i}$, we could consider the example illustrated in Fig. 2. The user request, ”I want a historical tour that begins at a cafe, avoiding crowds,” can be broken down into two subrequests. The first specifies the starting point at a cafe, which is a subrequest related to a POI, making it a ”start” ‘type’ with ‘pos’ as ”a cafe”, ‘neg’ as null, and ‘mustsee’ as true, indicating it’s an must visit POI. The second subrequest seeks a historical route that avoids crowds. For this, the ‘type’ is ”itinerary”, ‘pos’ is ”historical tour”, ‘neg’ is ”crowds”, and mustsee is false, indicating a preference rather than a constraint.
For those subrequests identified with the identified ‘type’ as ”POI” and the value of the key ‘mustsee’ is true, we leverage their ‘pos’ strings to conduct fuzzy string match over the POI names stored in the user-owned POI database to retrieve the must-see POIs $\small\mathcal{P}^{\operatorname{must}}=\{p^{\operatorname{must}}_{j}\}_{j=1}^{|\mathcal{P}^{\operatorname{must}}|}$. We provide the full prompt $\small\mathbb{P}_{RD}$ is available in Appendix F.1.
4.4. Preference-aware POI Retrieval
To generate itineraries that meet user requests, the initial step involves selecting appropriate POIs. Although LLMs possess advanced intelligence in language comprehension, their capabilities are limited by their context window size (Liu et al., 2023; Pal et al., 2023), and the cost is proportional to the input tokens.. Given the vast amount of POI data in the real world and the limitations on LLM inference speed that could hinder our system’s response time, we have adopted a embedding-based retrieval approach in this module.
Specifically, consider a single subrequest $\small\mathbf{r}_{i}$, we first employ an embedding model $\small\operatorname{E}_{\theta^{\prime}}$ to encode both the ‘pos’ and ‘neg’ fields:
(3) $\small\mathbf{e}^{\operatorname{pos}}_{i}=\operatorname{E}_{\theta^{\prime}}\left(\mathbf{r}_{i}^{\operatorname{pos}}\right);\,\mathbf{e}^{\operatorname{neg}}_{i}=\operatorname{E}_{\theta^{\prime}}\left(\mathbf{r}_{i}^{\operatorname{neg}}\right),$
where $\small\theta^{\prime}$ denotes the parameters of the embedding model $\small\operatorname{E}$, and $\small\mathbf{e}^{\operatorname{neg}}_{i},\mathbf{e}^{\operatorname{pos}}_{i}\in\mathbb{R}^{d}$ are two resulted $\small d$-dimensional embeddings.

### v4

# Yihong Tang1,2*, Zhaokai Wang3*, Ao Qu1,4*, Yihao Yan1*, Zhaofeng Wu4 Dingyi Zhuang1,4, Jushi Kai3, Kebing Hou1, Xiaotong Guo4 Jinhua Zhao4🖂, Zhan Zhao2🖂, Wei Ma5🖂

@@
Abstract
Citywalk, a recently popular form of urban travel, requires genuine personalization and understanding of fine-grained requests compared to traditional itinerary planning. In this paper, we introduce the novel task of Open-domain Urban Itinerary Planning (OUIP), which generates personalized urban itineraries from user requests in natural language. We then present ItiNera, an OUIP system that integrates spatial optimization with large language models to provide customized urban itineraries based on user needs. This involves decomposing user requests, selecting candidate points of interest (POIs), ordering the POIs based on cluster-aware spatial optimization, and generating the itinerary. Experiments on real-world datasets and the performance of the deployed system demonstrate our system’s capacity to deliver personalized and spatially coherent itineraries compared to current solutions. Source codes of ItiNera are available at https://github.com/YihongT/ITINERA.
@@ Abstract
envname-P envname#1
[Uncaptioned image]ItiNera: Integrating Spatial Optimization with Large Language Models for Open-domain Urban Itinerary Planning
Yihong Tang1,2*, Zhaokai Wang3*, Ao Qu1,4*, Yihao Yan1*, Zhaofeng Wu4 Dingyi Zhuang1,4, Jushi Kai3, Kebing Hou1, Xiaotong Guo4 Jinhua Zhao4🖂, Zhan Zhao2🖂, Wei Ma5🖂 1Tutu AI  2University of Hong Kong  3Shanghai Jiao Tong University 4Massachusetts Institute of Technology  5The Hong Kong Polytechnic University yihongt@connect.hku.hk  {wangzhaokai,json.kai}@sjtu.edu.cn  {qua,zfw,dingyi,xtguo,jinhua}@mit.edu {yanyihao,houkebing}@tutu-ai.com  zhanzhao@hku.hk  wei.w.ma@polyu.edu.hk
00footnotetext: Equal contribution. 🖂Corresponding authors.
1 Introduction
As a novel form of urban travel, citywalk Germano (2023) invites travelers to wander through city streets and immerse themselves in local culture, offering a more dynamic, immersive, and fine-grained travel experience compared to traditional tourism. Planning a citywalk is a complex urban itinerary planning problem Halder et al. (2024), involving travel-related information gathering, POI selection, route mapping, and customization for diverse user needs. Specifically, citywalk differs from traditional tourism by (1) Dynamic Information: involving rapidly changing POIs and needing up-to-date information on temporary events, (2) Personalization: prioritizing individual preferences over widely recognized POIs, and (3) Diverse Constraints: considering complex constraints like personal interests and accessibility requirements. An example of the OUIP problem is shown in Fig. 1.
Existing itinerary planning studies focus on traditional tourism. They consider coarse-grained user requirements such as geographical constraints Rani et al. (2018) and time budgets Hsueh and Huang (2019) to improve the quality of an itinerary Chen et al. (2013); Sylejmani et al. (2017). While these optimization-based approaches maintain the quality of POIs and spatial coherency, they struggle to address dynamic and detailed personal demands, leading to itineraries that lack personalization and diversity.
Recently, large language models (LLMs) OpenAI (2023) have shown impressive applications in understanding user needs and following instructions. However, their limitations in itinerary planning are evident Xie et al. (2024): (1) Pure LLMs cannot refer to specific POI lists, resulting in outdated or hallucinated POIs. (2) LLMs lack the optimization capabilities required for planning tasks, leading to suboptimal itineraries. Consequently, LLM-generated itineraries can be circuitous, lack detail, and include impractical information.
Figure 1: The OUIP problem and the OUIP system.
To address these limitations, in this work, we first define the Open-domain Urban Itinerary Planning (OUIP) problem, which involves generating personalized travel itineraries based on user requests in natural language. Then, we propose ItiNera, a holistic OUIP system that integrates spatial optimization with LLMs. ItiNera comprises five LLM-assisted modules: User-owned POI Database Construction (UPC), Request Decomposition (RD), Preference-aware POI Retrieval (PPR), Cluster-aware Spatial Optimization (CSO), and Itinerary Generation (IG), to deliver personalized and spatially coherent itineraries.
Our overall contributions are:
  • We introduce the OUIP problem to provide personalized urban travel itineraries based on users’ natural language inputs and propose metrics to measure the quality of generated itineraries.
  • We develop ItiNera, an LLM-based OUIP system that combines spatial optimization with LLMs to create fine-grained urban itineraries tailored to users’ requests.
  • Extensive experiments on the real-world dataset and performance of the deployed system show that ItiNera creates personalized, spatially coherent, and high-quality urban itineraries that meet user requirements.
Figure 2: An overview of the proposed ItiNera system.
2 Related Work
LLMs in Urban Applications
Since ChatGPT, LLMs have demonstrated strong knowledge and reasoning capabilities. Recent studies highlight the potential of LLMs in urban data processing Yan et al. (2023) and urban planning Zhou et al. (2024). These works reveal LLMs’ capabilities in predicting human mobility patterns Mo et al. (2023); Xue et al. (2022) and emphasize their predictive strength Wang et al. (2023). In transportation, LLMs contribute to traffic safety analysis Zheng et al. (2023a), enhance traffic forecasting de Zarzà et al. (2023), and automate accident report generation Zheng et al. (2023b) , showing their applicability in urban transportation. Leveraging LLMs for travel planning has recently gained public interest. TravelPlanner Xie et al. (2024) proposes a sandbox environment with various tools for benchmarking LLMs on multi-day travel planning, revealing LLMs’ current limitations for complex planning tasks. Unlike TravelPlanner, our system focuses on fine-grained OUIP, addressing urban itinerary planning within a single day, but can be seamlessly extended to multi-day travel planning.
Itinerary Planning (IP)
Current research on IP focuses on creating itineraries based on a set of POIs. Some methods directly optimize the spatial utilities of the itinerary, while others define IP as an Orienteering Problem (OP) and consider constraints that include time Zhang and Tang (2018); Hsueh and Huang (2019), space Rani et al. (2018), must-see POIs Taylor et al. (2018), categories Bolzoni et al. (2014) , and their combinations Gionis et al. (2014); Yochum et al. (2020), to indirectly ensure the spatial coherence and quality of the itinerary. However, their ability to personalize is limited. Recommendation-based methods Ho and Lim (2022); Tang et al. (2022) could be applied to the IP task, but they depend on historical user behavior data. Overall, existing IP methods struggle with open-domain, user natural-language inputs, failing to generate personalized itineraries, making them unsuitable for OUIP.
3 Methodology
We formalize the OUIP problem and explain how ItiNera generates itineraries, as shown in Fig. 2.
3.1 Open-domain Urban Itinerary Planning (OUIP) Problem
To enable personalized OUIP, an open-domain system is essential. Such a system allows users to freely express their diverse requirements and expectations, enabling the planning of urban itineraries tailored to their specific needs and purposes.
Formally, given a user request $r$ in natural language and the user-owned POI database $\mathcal{P}=\{p_{j}\}_{j=1}^{N}$, the OUIP problem aims to find an itinerary generator $\mathcal{G}$ to select and order a subset of POIs from $\mathcal{P}$ to create a coherent travel itinerary $I$ as an ordered list of POIs that optimally aligns with the user’s requests $r$ while adhering to spatio-temporal considerations: $I\sim\mathcal{G}\left(\mathcal{P}|r\right)$.
3.2 User-owned POI Database Construction
Travelers often have specific places they want to visit or particular requirements for the POIs in their itinerary. To ensure a fully personalized itinerary planning process, we have developed an automated pipeline that extracts POIs and relevant details from social media, catering to these individual needs. Users can input travel post links, and the pipeline uses LLMs to extract POIs and their descriptions, calls Map APIs and embedding models to obtain their locations and embeddings, and integrates the information into the user-owned POI database $\mathcal{P}$ and embedding database $\mathcal{E}$.
The user-owned POIs enable users to create personalized POI databases, maintain timely POI information, and customize travel itineraries, enhancing itinerary experiences. We execute a daily routine to aggregate POIs from trending posts across multiple cities and maintain an up-to-date, dynamic and comprehensive POI database. This database serves as the initial set of POIs for any new user, substantially mitigating the potential cold start issue for POI acquisition. The pipeline and the database format are detailed in §E and §B.
3.3 Request Decomposition
Upon receiving user requests, we use LLMs to structure and extract information. A single user request $r$ can be decomposed into multiple independent subrequests, each reflecting preferences at different levels and classified by granularity, specificity, and attitude. Granularity includes (1) POI-level and (2) itinerary-level subrequests. Specificity has (1) specific and (2) vague subrequests. Attitude distinguishes (1) positive subrequests (likes) and (2) negative subrequests (dislikes).
We prompt the LLM to decompose the user request $r$ based on these categories. Formally, we obtain the resulting set of structured subrequests $\mathcal{R}=\{\mathbf{r}_{i}\}_{i=1}^{|\mathcal{R}|}$ through: $\mathcal{R}\sim\operatorname{LLM}\left(\mathbb{P}_{RD}(r)\right)$, where $\mathbb{P}_{RD}$ wraps the request $r$ with instructions and examples (see §F.1). Here, ‘pos’ and ‘neg’ indicate attitude-level subrequests. ‘Mustsee’ is a boolean for specificity-level subrequests, and ‘type’ indicates granularity-level subrequests, which can be one of ["start" (POI-level origin), "end" (POI-level destination), "POI" (POI-level), "itinerary" (itinerary-level)].
3.4 Preference-aware POI Retrieval
After obtaining decomposed subrequests, we select POIs from the user-owned POI database $\mathcal{P}$ that match their preferences. While LLMs excel in language comprehension, they are limited by context window size and input token cost. Given the vast amount of POI data and LLM inference speed limitations, we design a preference-aware embedding-based retrieval approach. For a subrequest $\mathbf{r}_{i}$, we first use an embedding model $\operatorname{E}_{\theta^{\prime}}$ to encode the ‘pos’ and ‘neg’ fields: $\mathbf{e}^{\operatorname{pos}}_{i}=\operatorname{E}_{\theta^{\prime}}\left(\mathbf{r}_{i}^{\operatorname{pos}}\right);\,\mathbf{e}^{\operatorname{neg}}_{i}=\operatorname{E}_{\theta^{\prime}}\left(\mathbf{r}_{i}^{\operatorname{neg}}\right)$, where $\theta^{\prime}$ denotes the parameters of the $\operatorname{E}$, and $\mathbf{e}^{\operatorname{neg}}_{i},\mathbf{e}^{\operatorname{pos}}_{i}\in\mathbb{R}^{d}$ are embeddings.